OpenStudy (anonymous):

Currently, Sam's coach is three times older than Sam. When Sam started soccer 15 years ago, his coach was six times his age. How old is Sam now?

OpenStudy (anonymous):

Can set this up with variables x and y. X = Sam's Age, Y = Coach's Age. Our first equation: 3X = Y second equation: 6(X-15) = Y - 15, Let us expand this equation: 6X - 90 = Y - 15, or 6X - 75 = Y. Now let us set these equations equal to each other: 6X - 75 = Y and 3X = Y, so 6X - 75 = 3X 3X = 75, X = 25 - so at current age Sam is 25 years old and his Coach is 75 years old.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

You can double check by subtracting 15 from both ages and seeing if Sam's age (10) multiplied by 6 is equal to the Coachs age (minus 15) - which it is: 60. Sam is 25 currently not 30. People seem to forget that 15 years ago Sam's coach was also 15 years younger.